Daniel Mathews a footballer who is based in Sweden received a fine of $2 500 and a two-match suspension for use of performance-enhancing drugs. Daniel receives a match fee of $5 000 per match from the Sweden Footballers Board, plus additional appearance money of $10 000 per match from Adidas. He also pays $500 per match for comprehensive insurance against injury. In opportunity cost terms, the cost of this incident for Daniel Mathews is:


1
Expert's answer
2021-12-01T10:45:40-0500

Fine =$2500 and 2 match suspension

Let us call each match x


Hence the income for Mathew is

"I=5000x+10000x-500x"


For 2 match suspension Mathew will incur a cost of


"I=5000(2)+10000(2)-500(2)\\=\$29000"


the cost of this incident for Daniel Mathews is:


"=29000+2500"


=$31500